OpenStudy (anonymous):

4(x+6)/3=2x-1 solve for x

OpenStudy (anonymous):

4x+6 = 6x-3 9 = 2x x = 9/2

OpenStudy (anonymous):

negative good budy program says its wrong i see what you did though but theres no way of typing it out on the comp think of it as four next to x over six